let PS be the median of the triangle with vertices P(2,2),Q(6,-1)and R(7,3).the equation of the line passing through (1,-1) and parallel to PS is

Grade:12

2 Answers

Rinkoo Gupta
9 years ago
coordinate of S=(13/2,1)
slope of ps=(1-2)/13/2 -2)=(-1)/(9/2)=-2/9
Eq. of the line passing through (-1,1) and parallel to PS is
y-1=(-2/9)(x+1)
9(y-1)=-2x-2
9y-9+2x+2=0
2x+9y-7=0

Thanks & Regards
Rinkoo Gupta
AskIITians Faculty
Ankan Paul
6 years ago
coordinate of S=(13/2,1)slope of ps=(1-2)/13/2 -2)=(-1)/(9/2)=-2/9Eq. of the line passing through (-1,1) and parallel to PS isy+1=(-2/9)(x-1)9(y+1)=-2x+29y+9+2x-2=02x+9y+7=0A little correction but a big difference.